OpenStudy (anonymous):

how would i state the domain for this: f(x) = - 2 cot 3x

OpenStudy (anonymous):

is that a 3x or cot^3 (x)

OpenStudy (anonymous):

3x

OpenStudy (rock_mit182):

let's say U = 3x. Y = -2 cot (U). Do you remember the identity of cot (U) =

OpenStudy (rock_mit182):

?

OpenStudy (rock_mit182):

@kirarittgers

OpenStudy (anonymous):

no i dont know what that means

OpenStudy (rock_mit182):

cot (U) = \[\frac{ \cos (u) }{ \sin(u) }\]

OpenStudy (anonymous):

oh yeah i knew that

OpenStudy (anonymous):

but what do i do with that

OpenStudy (rock_mit182):

;) now, to find out your domain we have to avoid \[\frac{ 0 }{ 0 }\]

OpenStudy (rock_mit182):

cause \[\frac{ 0 }{ 0 }\] doesn't make sense in math is jut not allowed

OpenStudy (anonymous):

no shh i really knew that i just didnt know what you were asking and i still know that what i dont understan is how to state it

OpenStudy (rock_mit182):

so we need to avoid those values tha makes: \[\frac{ \cos u }{ \sin u }\]

OpenStudy (anonymous):

he gave you the identity, now he's trying to avoid sin U where sin U = 0 so that the function is undefined

OpenStudy (rock_mit182):

0 over o

OpenStudy (anonymous):

like i know it cant equal pi/3 or any other vertical asymtope but i dont know how to write that

OpenStudy (anonymous):

i know the values it cant equal though i need to know how to write it

OpenStudy (rock_mit182):

well it is getting harder to explain LOL

OpenStudy (rock_mit182):

Do you have how to graph that function ?

OpenStudy (anonymous):

how bout you just tell me the answer lol because i already know everything i dont need explaining i need to know how its supposed to be writen

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

ive already graphed it

OpenStudy (anonymous):

do you understand what im asking like you know how you write it in parenthesis i need to know how im supposed to put the values in there

OpenStudy (rock_mit182):

of course i do , but what i cannot do, here, is just give ramdom answers

OpenStudy (anonymous):

its not random but oh well